y=2x+1 find the slope and y-intercept
y=-3x-2 find the slope and y-intercept
y=-3/4x+1 find the slope and y-intercept

Found 2 solutions by tutorcecilia, bucky:
Answer by tutorcecilia(2152)

Slopes and y-intercepts:
Compare each equation with (y=mx+b)
y=any y-value
m=the slope of the line
x=any corresponding x-value
b=the y-intercept
.
y=2x+1 find the slope and y-intercept
y=mx+b
Slope=m=2
y-intercept=b=1
.
y=-3x-2 find the slope and y-intercept
y=mx+b
Slope=m=-3
y=intercept = b=-2
.
y=-3/4x+1 find the slope and y-intercept
y=mx+b
Slope=m=-3/4
y=intercpet=b=1

Answer by bucky(2189)

The equations that you are given are all in the slope-intercept form of:
.
y = mx + b
.
in which m is the slope and b is the y-intercept. Therefore, you can identify the slope and
the y-intercept by inspection. The slope is the number that multiplies the x in the
equation and the number that follows the x term is the y-intercept.
.
Using these two rules, you can tell that in the equation:
.
y = 2x + 1
.
the slope (multiplier of x) is +2 and the y-intercept is +1 (the number that follows the
term that contains the x).
.
Similarly, in the second equation:
.
y = -3x - 2
.
the slope is -3 and the y-intercept is -2
.
And in the final equation:
.
y = -(3/4)x + 1
.
the slope is -3/4 and the y-intercept is + 1
.
Hope this helps you to understand the problem